OpenStudy (love_to_love_you):

Please help! Will do anything. Write an equation for each line in point-slope form and then convert it to standard form. Show steps. 1. through (2, 3) and (3, 5)

OpenStudy (love_to_love_you):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Whats the slope of this line?

OpenStudy (love_to_love_you):

I think I have to figure it out

jimthompson5910 (jim_thompson5910):

yes you would use the slope formula

OpenStudy (love_to_love_you):

5-3/3-2 2/1 slope=2

OpenStudy (love_to_love_you):

y - y1 = m(x-x1) y - (3) = 2(x - 2) y-3=2x-4 Im stuck here

jimthompson5910 (jim_thompson5910):

y - 3 = 2(x - 2) is the equation in point-slope form

jimthompson5910 (jim_thompson5910):

solve for y y - 3 = 2(x - 2) y - 3 = 2x - 4 y = 2x - 4+3 y = 2x - 1 the equation is now in slope intercept form

OpenStudy (love_to_love_you):

I need standard form, is that standard form?

jimthompson5910 (jim_thompson5910):

to get it into standard form, you need it in Ax+By = C form

jimthompson5910 (jim_thompson5910):

usually A > 0

jimthompson5910 (jim_thompson5910):

y = 2x - 1 y+1 = 2x 1 = 2x - y 2x - y = 1 and it's now in standard form

OpenStudy (love_to_love_you):

in the first one you helped me with was it standard or slope that you gave me?

jimthompson5910 (jim_thompson5910):

I don't remember

jimthompson5910 (jim_thompson5910):

You'll have to post it

OpenStudy (love_to_love_you):

I figured it out yes it was. Could you please help me with another? Sorry its just I am completely screwed because I have a horrible teacher and I have to turn in a study guide and I have like 13 questions I need help with. You don't need to help me with them (unless you want to), but I would appreciate it if you could!!

OpenStudy (love_to_love_you):

there was 55 questions in total I just need help with these.

jimthompson5910 (jim_thompson5910):

I won't help with all 13, but I'll help with a few more.

OpenStudy (love_to_love_you):

thank you!!

OpenStudy (love_to_love_you):

Draw a scatter plot of each set of data. Decide whether a linear model is reasonable. If so, describe the correlation. Then draw a trend line and write it's equation. Predict the value of y when x is 15. 1. {(3, 5), (4, 7), (5, 9), (7, 10), (8, 10), (9, 11), (10, 13)} 2. x- 0, 3, 6, 9, 12 y- 17.5, 35.4, 50.5, 60.6, 66.3 this is a table ^

jimthompson5910 (jim_thompson5910):

what kind of calculator do you have?

OpenStudy (love_to_love_you):

it broke

jimthompson5910 (jim_thompson5910):

do you have microsoft excel?